A.Calculate the resultant of following vectors 4.0 m/s and 10 .0 m/s, with an angle of 90 degrees between them.

B. Calculate the resultant of the following vectors 4.0 N and 10 .0 N, with an angle of 0 degrees between them.
Answer #1

A)

Resultant=sqrt(A^2+B^2+2ABcos(theta))

So here theta=90 degree

Resultant=sqrt(16+100)=10.8 N

B

Here theta=0

Resultant=14 N